OpenStudy (anonymous):

find the distance between the points (-1,3) and (5,0)

OpenStudy (anonymous):

d= (-1-5) ^2 + (3-0) ^2 d= (-6)^2 + (3)^2 d= 36 + 9 d= 45 d= 3 root 5.